# On the inadequacy of nudging data assimilation algorithms for non-dissipative systems

### v2

@@ 4.2. Computational Results
In this section we will detail the numerical methods we utilize to simulate Sections˜4.1 and 4.1. All simulations depicted in this study were performed Matlab (R2023a) using our own code. For this computational study we implemented the incompressible Euler equations on the 2D periodic torus $\mathbb{T}^{2}=[0,2\pi)\times[0,2\pi)$, subject to periodic B.C. with period with pseudo-spectral methods. We implement the equations by using the explicit Euler time-stepping scheme with the nonlinear computed explicitly. In the simulations in this chapter we utilize pseudo-spectral methods (see e.g., [14, 72, 75] for textbook descriptions of these methods), that is, derivatives were computed using Matlab’s implementation of the n-dimensional fast Fourier transform (fftn, which relies on FFTW). We utilized the standard 2/3 dealiasing rule to compute the nonlinear term, that is, we zeroed out the modes above wavenumber $N/3$, where $N$ is number of gridpoints in one direction of the periodic box. Moreover, to reduce the number of FFT operations, the Basdevant formula was used (see [30] for further discussion). We utilize the spatial resolution $N=2^{8}$ and time step $\Delta t=0.001$.
We remark that while our analytical results illustrate the failure of nudging to recover the true solution, effective testing of the AOT algorithm in recovering solutions to Section˜4.1 remains problematic. This is due to the fact that the AOT algorithm is typically shown to recover dissipative dynamical systems only asymptotically in time. Thus if one wishes to faithfully demonstrate the effectiveness of the AOT algorithm it requires running simulations of Section˜4.1 for long amounts of time. This in itself is problematic, as it was shown in [29] that solutions to Section˜4.1 supported on a finite number of Fourier modes for all times $t\in[0,\infty)$ must necessarily be constant in time. This indicates that solutions that are evolving in time develop arbitrary fine length scales and thus one cannot accurately simulate the long time dynamics using a fixed spatial resolution. However, we note that steady-state solutions are in fact enough to illustrate the failing of the AOT algorithm on the incompressible Euler equations.
For our reference solution we utilize the Taylor-Green vortex given with frequency $k=15$ and $c=10^{-4}$, given as follows
@@
We see in Figure˜11(a) that observing all Fourier modes $\boldsymbol{k}$ with $\mathinner{\!\left\lvert\boldsymbol{k}\right\rvert}\leq 21$ is insufficient to recover the true solution. We note that in the spectral plots Figures˜11(b) and 12(b) the vertical lines indicate the observational frequency and dealiasing cutoff frequencies, given in blue and red, respectively. While $M=21$ is insufficient to recover the true solution, we find that $M=22$ is sufficient, which is of course unsurprising as this is exactly the regime in which the entire solution is in fact observed. When $M$ is chosen smaller than the frequency utilized in the construction of the Taylor-Green vortex the solution $\boldsymbol{v}(\boldsymbol{x},t)\equiv\boldsymbol{0}$ holds for all time $t\in[0,1]$.
@@
Figure 11. Nudging with 2121 observed modes fails to recover reference solution with nudging with $\mu=100$. The nudged solution is initialized to be identically zero and the reference solution is initialized as in Equation˜41 with $k=15$. Note that $M=21$ gives the radius of the observed frequencies, and the effective frequency for $k=15$ is in fact $\mathinner{\!\left\lvert(15,15)\right\rvert}=\sqrt{2}(15)\approx 21.21.$
@@
Figure 12. Nudging with $22$ observed modes fails to recover reference solution with nudging with $\mu=100$. The nudged solution is initialized to be identically zero and the reference solution is initialized as in Equation˜41 with $k=15$. Note that $M=22$ gives the radius of the observed frequencies, and the effective frequency for $k=15$ is in fact $\mathinner{\!\left\lvert(15,15)\right\rvert}=\sqrt{2}(15)\approx 21.21.$
@@ 5. Conclusion
Based on the results of the simulation presented in Section˜3.2 and Section˜4.2, we conclude that the AOT data assimilation algorithm is unsuitable for non-dissipative or partially dissipative dynamical systems, such as the incompressible Euler and KdV equations, Sections˜4.1 and 19, respectively, and the partially dissipative Lorenz system, Equation˜1. In fact, we conclude that the pathologic counterexamples used in Sections˜3.2.1 and 4 present an insurmountable challenge for any method of data assimilation to recover. In particular, we note that counterexample we utilize in Figure˜4 will fail for any method of data assimilation, as the true solution has no discernible impact on the observed modes, and so it is impossible to reliably tell the true solution apart from the zero solution (or one of an infinite number of solutions one can readily construct) based solely on the observations. Moreover, the counter examples we considered in this work are not specific to the KdV or Euler equations. The particular solutions considered in Proposition˜3.5 were constructed due to the periodic invariance induced by the convolutional Fourier mode structure of the nonlinear term (see Corollary˜3.4), and thus one can construct similar counter examples for other non-dissipative systems with a similar nonlinearity.
In addition to showing that data assimilation simply doesn’t work for certain pathological counterexamples, we note that the extent to which data assimilation does work for the classical KdV equation may be highly dependent on the choice of initial data for the nudged system, $v^{in}$. Indeed, while it was proved in [4] that for 2D NSE the choice of initial data $v^{in}$ was arbitrary, we have seen here that data assimilation will simply fail if not initialized properly, as it does in Figure˜4. The problem of initialization is one of the core strengths of the nudging algorithm, and so it is worth noting that this property fails to hold for a system without finitely many determining modes.
We note that data assimilation could be made to work if one were to assume more of the true solution. For instance, if one were to assume that the solution was supported on an annulus in Fourier space, one could eliminate all the examples that we study here. We caution anyone against doing this, however, as imposing these sorts of restrictions on the solution can trivialize the dynamics of equations in a way that may not be physical.

### v3

@@ 4.2. Computational Results
In this section we will detail the numerical methods we utilize to simulate Equations˜34 and 35. For this computational study we implemented the incompressible Euler equations on the 2D periodic torus $\mathbb{T}^{2}=[0,2\pi)\times[0,2\pi)$, subject to periodic B.C. with period $2\pi$ using pseudo-spectral methods. We implement the equations by using the explicit Euler time-stepping scheme with the nonlinear computed explicitly. In the simulations in this chapter we utilize pseudo-spectral methods (see e.g., [14, 72, 75] for textbook descriptions of these methods), that is, derivatives were computed using Matlab’s implementation of the n-dimensional fast Fourier transform (fftn, which relies on FFTW). We used the standard 2/3 dealiasing rule to compute the nonlinear term, that is, we zeroed out the modes above wavenumber $N/3$, where $N$ is number of gridpoints in one direction of the periodic box. Moreover, to reduce the number of FFT operations, the Basdevant formula was used (see [30] for further discussion). We utilize the spatial resolution $N=2^{8}$ and time step $\Delta t=0.001$.
We remark that while our analytical results illustrate the failure of nudging to recover the true solution, effective testing of the AOT algorithm in recovering solutions to Equation˜34 remains problematic. This is due to the fact that the AOT algorithm is typically shown to recover dissipative dynamical systems only asymptotically in time. Thus if one wishes to faithfully demonstrate the effectiveness of the AOT algorithm it requires running simulations of Equation˜34 for long amounts of time. This in itself is problematic, as it was shown in [29] that solutions to Equation˜34 supported on a finite number of Fourier modes for all times $t\in[0,\infty)$ must necessarily be constant in time. This result directly shows that Fourier-based pseudo-spectral schemes cannot adequately approximate non-steady state solutions to Equation˜34 over arbitrarily long time intervals as they necessarily need to resolve the developing length scales finer than the fixed spatial resolution. This issue could potentially be alleviated by using methods with adaptive mesh refinement, however at some point the computer running the simulation will run out of the resources necessarily to adequately capture the developing fine length scales. However, steady-state solutions are in fact enough to illustrate the failing of the AOT algorithm on the incompressible Euler equations.
For our reference solution we utilize the Taylor-Green vortex given with frequency $k=15$ and $c=10^{-4}$, given as follows (see also Figure˜10)
@@
We see in Figure˜11(a) that observing all Fourier modes $\boldsymbol{k}$ with $\mathinner{\!\left\lvert\boldsymbol{k}\right\rvert}\leq 21$ is insufficient to recover the true solution. In the spectral plots Figures˜11(b) and 12(b) the vertical lines indicate the observational frequency and dealiasing cutoff frequencies, given in blue and red, respectively. While $M=21$ is insufficient to recover the true solution, we find that $M=22$ is sufficient, which is of course not surprising as this is exactly the regime in which the entire solution is in fact observed. When $M$ is chosen smaller than the frequency used in the construction of the Taylor-Green vortex the solution $\boldsymbol{v}(\boldsymbol{x},t)\equiv\boldsymbol{0}$ holds for all time $t\in[0,1]$.
@@
Figure 11. Nudging with 2121 observed modes fails to recover reference solution with nudging with $\mu=100$. The nudged solution is initialized to be identically zero and the reference solution is initialized as in Equation˜42 with $k=15$. Note that $M=21$ gives the radius of the observed frequencies, and the effective frequency for $k=15$ is in fact $\mathinner{\!\left\lvert(15,15)\right\rvert}=\sqrt{2}(15)\approx 21.21.$
@@
Figure 12. Nudging with $22$ observed modes fails to recover reference solution with nudging with $\mu=100$. The nudged solution is initialized to be identically zero and the reference solution is initialized as in Equation˜42 with $k=15$. Note that $M=22$ gives the radius of the observed frequencies, and the effective frequency for $k=15$ is in fact $\mathinner{\!\left\lvert(15,15)\right\rvert}=\sqrt{2}(15)\approx 21.21.$
@@ 5. Conclusion
Based on the results of the simulation presented in Section˜3.2 and Section˜4.2, we conclude that the AOT data assimilation algorithm is unsuitable for recovering non-dissipative or partially dissipative dynamical systems, such as the incompressible Euler and KdV equations, Equations˜34 and 19, respectively, and the partially dissipative Lorenz system, Equation˜1. In fact, we conclude that the pathological counterexamples used in Sections˜3.2.1 and 4 present an insurmountable challenge for any method of data assimilation to recover. In particular, the counterexample we utilize in Figures˜4 and 11 will fail for any method of data assimilation, as the true solution has no discernible impact on the observed modes, and so it is impossible to reliably tell the true solution apart from the zero solution (or one of an infinite number of solutions one can readily construct) based solely on the observations. Moreover, similar counter examples can be readily constructed for other non-dissipative dynamical systems, as the counter examples we considered in this work are not specific to the KdV or Euler equations, but rather are a consequence of the invariance of certain periodic subspaces under the flow. The particular solutions considered in Propositions˜3.5 and 4.3 were constructed due to the periodic invariance induced by the convolutional Fourier mode structure of the nonlinear term (see Corollaries˜3.4 and 4.2).
In addition to showing that data assimilation simply does not work for certain pathological counterexamples, we note that the extent to which data assimilation does work for the classical KdV equation may be highly dependent on the choice of initial data for the nudged system, $v^{in}$. Without dissipation in the dynamical system, the errors in the initial data may persist indefinitely, preventing the accurate recovery of the reference solution. Indeed, while it was proved in [4] that for 2D NSE the choice of initial data $v^{in}$ was arbitrary, we have seen here that data assimilation will simply fail if not initialized properly, as it does in Figure˜4. Overcoming the problem of initialization is one of the core strengths of the nudging algorithm, and so it is worth noting that this property fails to hold for a system without finitely many determining modes.
It is worth mentioning that that data assimilation could be made to work if one were to assume more of the true solution. For instance, if one were to assume that the solution was supported on an annulus in Fourier space, one could eliminate all the examples that we study here. We caution against doing this, however, as imposing these sorts of restrictions on the solution can trivialize the dynamics of equations in a way that may not be physical.